# The Mirage of Performance Gains: Why Contrastive Decoding Fails to Address Multimodal Hallucination

### v2

@@ Abstract
Contrastive decoding strategies are widely used to reduce hallucinations in multimodal large language models (MLLMs). These methods work by constructing contrastive samples to induce hallucinations and then suppressing them in the output distribution. However, this paper demonstrates that such approaches fail to effectively mitigate the hallucination problem. The performance improvements observed on POPE Benchmark are largely driven by two misleading factors: (1) crude, unidirectional adjustments to the model’s output distribution and (2) the adaptive plausibility constraint, which reduces the sampling strategy to greedy search. To further illustrate these issues, we introduce a series of spurious improvement methods and evaluate their performance against contrastive decoding techniques. Experimental results reveal that the observed performance gains in contrastive decoding are entirely unrelated to its intended goal of mitigating hallucinations. Our findings challenge common assumptions about the effectiveness of contrastive decoding strategies and pave the way for developing genuinely effective solutions to hallucinations in MLLMs.
@@ 1 Introduction
The hallucination problem (Li et al., 2023b; Liu et al., 2023a; Lovenia et al., 2023; Gunjal et al., 2023) in multimodal large language models (MLLMs) (Bavishi et al., 2023; Chen et al., 2024a; Ye et al., 2023; Chen et al., 2023a) refers to the generation of outputs that are factually incorrect or misaligned with the input data. This issue arises from challenges in aligning diverse data modalities, such as text and images, which amplify reasoning errors. Such hallucinations can have serious consequences in critical domains, including autonomous driving (Mai et al., 2023; Liu et al., 2023c; Chen et al., 2023b; Wu et al., 2023) (e.g., false object detection leading to accidents) and healthcare (Wang et al., 2023; Hu et al., 2023) (e.g., incorrect diagnostic interpretations).
@@
Contrastive decoding methods (Wu et al., 2022; Gupta et al., 2022; Niu et al., 2021) are widely recognized as an effective approach to addressing object hallucination in generative models. As illustrated in Figure 1, these methods construct contrastive samples designed to induce hallucinations, then suppress the corresponding output distributions, ensuring closer alignment between model outputs and visual inputs. Representative approaches within this framework include Visual Contrastive Decoding (VCD) (Leng et al., 2023), Instruction Contrastive Decoding (ICD) (Wang et al., 2024b), and Self-Introspective Decoding (SID) (Huo et al., 2024). Their training-free nature and purported ability to address hallucinations have made them highly regarded in the field.
@@
Although methods like VCD have demonstrated remarkable performance improvements on the POPE benchmark (Li et al., 2023c), we reveal in Section 4 that these results are highly misleading. In reality, these methods fail to effectively address model hallucination. The observed performance gains on the POPE benchmark are primarily driven by two factors:
@@
To expose the misleading nature of the performance improvement in the first scenario, we implemented two forced distribution adjustment algorithms in Section 5.1 to show that the apparent gains of contrastive decoding on the POPE Benchmark are not genuine. The methods are as follows: (1) Prompt-Based Adjustment, where we added a hint to the instruction, such as ”Whenever possible, please select Yes,” to bias outputs toward ”Yes”; and (2) Output Layer Modification, where we altered the output layer to favor ”Yes” when the probabilities for ”Yes” and ”No” were similar. Although neither method mitigates hallucinations, both achieved performance gains comparable to those of contrastive decoding, confirming that these improvements do not represent a genuine solution to the problem.
@@
To highlight the misleading nature of the performance improvement in the second scenario, we incorporated the adaptive plausibility constraint into the standard sampling strategy and compared its predictions with those from contrastive decoding in Section 5.2. The experimental results reveal that, despite having no theoretical connection to hallucination mitigation, the adaptive plausibility constraint accounts for nearly all the performance gains attributed to contrastive decoding. This finding underscores that the contrastive decoding methods, in essence, fail to mitigate hallucinations.
@@ 2.1 Multimodal Large Language Models
The evolution of MLLMs (Chen et al., 2024b, c) has progressed from BERT-based decoders (Li et al., 2020, 2021) to advanced LLM architectures (Touvron et al., 2023a, b; Meta, 2024), enabling more effective multimodal relationship modeling (Chiang & Li, 2023; Taori et al., 2023; Bai et al., 2023a). Models such as BLIP-2 (Li et al., 2023a) and MiniGPT-4 (Zhu et al., 2023) employ Q-Former mechanisms to enhance the alignment between visual and textual inputs, facilitating more precise cross-modal interactions. InstructBLIP (Dai & et al, 2023) extends this framework by integrating task-specific instructions, improving the model’s ability to interpret context-sensitive visual semantics. Meanwhile, LLaVA (Liu et al., 2023b, 2024; Li et al., 2024) and Qwen-VL (Bai et al., 2023b) adopt simpler linear projection methods that streamline alignment, leading to superior performance in vision-language tasks. Despite these advancements, hallucination remains a persistent challenge that warrants further investigation.
2.2 Contrastive Decoding Strategies
@@
Contrastive decoding (Yan et al., 2023; Zhibo et al., 2023; Han et al., 2022) are widely recognized as effective in addressing object hallucination in generative models. Visual Contrastive Decoding (VCD) (Leng et al., 2023) addresses object hallucination by comparing output distributions generated from standard visual inputs and distorted visual inputs. This approach reduces the model’s dependence on linguistic priors within integrated LLMs and minimizes the impact of statistical biases in MLLM pretraining corpus. Instruction Contrastive Decoding (ICD) (Wang et al., 2024b), in contrast, focuses on the role of instruction perturbations in amplifying hallucinations. By examining the differences in output distributions between standard and perturbed instructions, ICD detects hallucination-prone content and mitigates its impact effectively.
Building upon these two hallucination mitigation methods, numerous approaches, including Adaptive Focal-Contrast Decoding (HALC) (Chen et al., 2024d), Self-Introspective Decoding (SID) (Huo et al., 2024), and Visual Layer Fusion Contrastive Decoding (VaLiD) (Wang et al., 2024a), have been developed based on similar principles. Although these methods have demonstrated substantial performance improvements on the POPE Benchmark, we will show that these improvements are, in fact, entirely unrelated to the original objective of hallucination mitigation.
3 Contrastive Decoding for Hallucinations

### v3

# The Mirage of Performance Gains: Why Contrastive Decoding Fails to Mitigate Object Hallucinations in MLLMs?

@@ Abstract
Contrastive decoding strategies are widely used to reduce object hallucinations in multimodal large language models (MLLMs). These methods work by constructing contrastive samples to induce hallucinations and then suppressing them in the output distribution. However, this paper demonstrates that such approaches fail to effectively mitigate the hallucination problem. The performance improvements observed on POPE Benchmark are largely driven by two misleading factors: (1) crude, unidirectional adjustments to the model’s output distribution and (2) the adaptive plausibility constraint, which reduces the sampling strategy to greedy search. To further illustrate these issues, we introduce a series of spurious improvement methods and evaluate their performance against contrastive decoding techniques. Experimental results reveal that the observed performance gains in contrastive decoding are entirely unrelated to its intended goal of mitigating hallucinations. Our findings challenge common assumptions about the effectiveness of contrastive decoding strategies and pave the way for developing genuinely effective solutions to hallucinations in MLLMs. The source code is available at https://github.com/ustc-hyin/cd_rethink
@@ 1 Introduction
The hallucination problem [1, 2, 3, 4] in multimodal large language models (MLLMs) [5, 6, 7] refers to the generation of outputs that are factually incorrect or misaligned with the input data. This issue arises from challenges in aligning diverse data modalities, such as text and images, which amplify reasoning errors. Such hallucinations can have serious consequences in critical domains, including autonomous driving [8, 9, 10, 11] (e.g., false object detection leading to accidents) and healthcare [12, 13] (e.g., incorrect diagnostic interpretations).
@@
Contrastive decoding methods [14, 15, 16] are widely recognized as an effective approach to addressing object hallucination in generative models. As illustrated in Figure˜1, these methods construct contrastive samples designed to induce hallucinations, then suppress the corresponding output distributions, ensuring closer alignment between model outputs and visual inputs. Representative approaches within this framework include Visual Contrastive Decoding (VCD) [17], Instruction Contrastive Decoding (ICD) [18], and Self-Introspective Decoding (SID) [19]. Their training-free nature and purported ability to address hallucinations have made them highly regarded in the field.
@@
Although methods like VCD have demonstrated remarkable performance improvements on the POPE benchmark [1], we reveal in Section˜4 that these results are highly misleading. In reality, these methods fail to effectively address model hallucination. The observed performance gains on the POPE benchmark are primarily driven by two factors:
@@
To expose the misleading nature of the improvement in the first scenario, we implemented two forced distribution adjustment algorithms in Section˜5.1 to show that the apparent gains of contrastive decoding on the POPE Benchmark are not genuine. The methods are as follows: (1) Prompt-Based Adjustment, where we added a prompt to the instruction, such as "Whenever possible, please select Yes." to bias outputs toward "Yes"; and (2) Output Layer Modification, where we altered the output layer to favor "Yes" when the probabilities for "Yes" and "No" were similar. Although neither method mitigates hallucinations, both achieved performance gains comparable to those of contrastive decoding, confirming that these improvements do not represent a genuine solution to the problem.
@@
To highlight the misleading nature of the performance improvement in the second scenario, we incorporated the adaptive plausibility constraint into the standard sampling strategy and compared its predictions with those from contrastive decoding in Section˜5.2. The experimental results reveal that, despite having no theoretical connection to hallucination mitigation, the adaptive plausibility constraint accounts for nearly all the performance gains attributed to contrastive decoding. This finding underscores that the contrastive decoding methods, in essence, fail to mitigate hallucinations.
@@ 2 Related Work
Multimodal Large Language Models. The evolution of MLLMs [20, 21] has progressed from BERT-based decoders [22, 23] to advanced LLM architectures [24, 25], enabling more effective multimodal relationship modeling [26, 27]. Models such as BLIP-2 [28] and MiniGPT-4 [29] employ Q-Former mechanisms to enhance the alignment between visual and textual inputs, facilitating more precise cross-modal interactions. InstructBLIP [30] extends this framework by integrating task-specific instructions, improving the model’s ability to interpret context-sensitive visual semantics. Meanwhile, LLaVA [31, 32] and Qwen-VL [33] adopt simpler linear projection methods that streamline alignment, leading to superior performance in vision-language tasks. Despite these advancements, hallucination remains a persistent challenge that warrants further investigation.
@@
Contrastive Decoding Strategies. Contrastive decoding [34, 35, 36] are widely recognized as effective in addressing object hallucination in generative models. Visual Contrastive Decoding (VCD) [17] addresses object hallucination by comparing output distributions generated from standard visual inputs and distorted visual inputs. This approach reduces the model’s dependence on linguistic priors within integrated LLMs and minimizes the impact of statistical biases in MLLM pretraining corpus. Instruction Contrastive Decoding (ICD) [18], in contrast, focuses on the role of instruction perturbations in amplifying hallucinations. By examining the differences in output distributions between standard and perturbed instructions, ICD detects hallucination-prone content and mitigates its impact effectively. Building upon these two hallucination mitigation methods, numerous approaches, including Adaptive Focal-Contrast Decoding (HALC) [37], Self-Introspective Decoding (SID) [19], and Visual Layer Fusion Contrastive Decoding (VaLiD) [38], have been developed based on similar principles. Although these methods have demonstrated substantial performance improvements on the POPE Benchmark, we will show that these improvements are, in fact, entirely unrelated to the original objective of hallucination mitigation.
3 Preliminary
This section outlines the main components and operational workflows of three leading hallucination mitigation methods: VCD, ICD, and SID, as illustrated in Figure˜1. All three adopt contrastive decoding strategies to reduce hallucinations and improve consistency with the visual input. The POPE Benchmark is also discussed as one of the most important metrics for evaluating their performance. Further technical details of VCD, ICD, and SID can be found in Appendix˜A.
Vanilla Decoding. We consider a MLLM parametrized by . The model takes as input a textual query and a visual input , where provides contextual visual information to assist the model in generating a relevant response to the textual query. The response is sampled auto-regressively from the probability distribution conditioned on the query and the visual context . Mathematically, this can be formulated as:
Visual Contrastive Decoding. VCD is a corrective strategy aimed at reducing hallucinations in MLLMs. Given a textual query and its corresponding image, the model produces two output distributions: one from the original image and another from a perturbed variant, typically generated by applying controlled modifications such as Gaussian noise. By evaluating the divergence between these distributions, VCD constructs a contrastive output distribution that improves the reliability and factual accuracy of the model’s response.
Instruction Contrastive Decoding. ICD addresses hallucinations by leveraging the observation that instruction perturbations, particularly those involving negative prefixes, increase uncertainty in multimodal alignment. ICD operates by first amplifying the probabilities of hallucinated concepts, then systematically detaching these from the original output distribution. This contrastive process reduces the model’s vulnerability to object hallucinations.
Self-Introspective Decoding. SID extends the ideas of VCD and ICD by addressing a key limitation: perturbing the full input can inject too much noise, making it harder to produce useful hallucinations. To resolve this, SID adaptively prunes the input by keeping only a small set of image tokens with low attention scores after the early decoder layers (see Figure˜1, far right). This focused adjustment encourages vision-and-text hallucinations to emerge during decoding. These hallucinated elements are then separated from the original output distribution to improve the model’s reliability.
Adaptive Plausibility Constraint. One key challenge inherent in the three aforementioned methods is the risk of indiscriminate penalization across the entire output space, which can unintentionally suppress valid predictions and, paradoxically, favor the generation of implausible tokens. To mitigate this, all three methods incorporate an adaptive plausibility constraint. This constraint dynamically adjusts penalization based on confidence scores derived from the model’s output distribution, conditioned on the original visual input . Formally, the constraint is defined as:
Here, represents the output vocabulary of the multimodal large language model (MLLM), and is a hyperparameter controlling the truncation threshold of the next-token distribution. A higher value of results in more aggressive truncation, thereby retaining only the most probable tokens.
Polling-based Object Probing Evaluation. POPE [1, 39] offers a robust framework for evaluating object hallucinations in multimodal large language models (MLLMs). Departing from caption-based methods, it frames hallucination detection as a binary task using direct Yes-or-No questions (e.g., “Is there a chair in the image?”), enabling clearer interpretation. The benchmark ensures a balanced distribution of “Yes” and “No” samples (50% each). Contrastive decoding–based mitigation strategies have demonstrated their effectiveness primarily through improved performance on POPE, reinforcing their credibility within the research community.
4 Misleading Performance Improvement
In this section, we highlight two misleading factors contributing to the performance improvement of contrastive decoding methods on the POPE Benchmark: (1) Unidirectional output adjustment skews the model towards generating more "Yes" outputs, leading to a balanced distribution in certain datasets. (2) The adaptive plausibility constraint degrades sampling decoding strategy into greedy search, resulting in deceptively improved outcomes.
| Dataset | COCO Random | GQA Adversarial | ||
| Method | Acc % | Yes % | Acc % | Yes % |
| Greedy | 87.1 | 39.2 | 80.9 | 54.0 |
| VCD | 88.6 | 46.4 | 78.0 | 63.3 |
| SID | 87.9 | 42.3 | 79.9 | 57.8 |
| Dataset | COCO Random | GQA Adversarial | ||
| Method | Acc % | Yes % | Acc % | Yes % |
| Greedy | 87.1 | 39.2 | 80.9 | 54.0 |
| VCD-C | 76.7 | 28.2 | 71.5 | 41.3 |
| SID-C | 79.0 | 23.6 | 74.2 | 43.1 |
4.1 Unidirectional Output Adjustment
In this subsection, we illustrate how contrastive decoding algorithms can deceptively enhance the performance of MLLMs on the POPE Benchmark by applying targeted, unidirectional modifications to the output distribution. We begin by evaluating the performance of various contrastive decoding methods on the MSCOCO [40] and GQA [41] datasets, analyzing both accuracy and the distribution of the model’s responses. For this study, we selected LLaVA-v1.5-7B as the foundational MLLM, using a greedy search decoding strategy.
As shown in Table˜1, both VCD and SID significantly skewed the model’s output distribution toward “Yes” across all subsets. On the MSCOCO-Random dataset, where the original output distribution was skewed toward "No," VCD and SID corrected this imbalance, resulting in a more balanced distribution and improved accuracy. Conversely, for the GQA-Adversarial subset, where the output distribution was already biased toward "Yes," these methods intensified the skew, ultimately reducing prediction accuracy.
We further illustrate how model outputs change after applying contrastive decoding methods, providing a clearer understanding of their performance improvements. As shown in Figure˜2, the method primarily alters predictions from "No" to "Yes," significantly outpacing the reverse. On the MSCOCO-Random dataset, where the output distribution is initially skewed toward "No," this adjustment converts many false negatives into true positives, thereby improving accuracy. Conversely, on the GQA-Adversarial dataset, which is biased toward "Yes," these modifications lead to the misclassification of numerous true negatives as false positives, resulting in a performance decline. For more details on prediction shifts, please refer to Appendix˜B.
To understand why contrastive decoding methods consistently increase the likelihood of a "Yes" response, we analyzed the output distribution generated from contrastive samples, as shown in Table˜2. In Section˜3, we proposed that the primary function of contrastive samples is to induce hallucinations, allowing contrastive decoding to subsequently filter out these hallucinated elements from the output distribution. However, the results in Table˜2 reveal that this objective was entirely unmet. Most outputs derived from contrastive samples were incorrect, not due to successfully induced hallucinations, but because the model overwhelmingly favored "No" responses. This severe bias in the output distribution led to a significant decline in accuracy. A more detailed explanation of why outputs generated from contrastive inputs are biased toward “No” is provided in Appendix˜C.
Based on the above discussion, the practical performance of contrastive decoding methods is illustrated in the lower section of Figure˜1. The output distribution derived from the contrastive inputs is heavily biased toward "No." However, the contrastive decoding method suppresses this content in the original output, thereby unilaterally increasing the model’s likelihood of answering "Yes." Whether the model’s performance on the dataset improves depends heavily on whether this increased "Yes" frequency leads to a more balanced output distribution. However, as shown in Figure˜3, the output distribution of MLLMs tends to be biased toward "No" in most data subsets. Therefore, contrastive decoding methods still manage to achieve a strong overall performance on the POPE Benchmark.
4.2 Sampling Decoding Degradation
Notably, the POPE Benchmark, which requires models to answer "Yes" or "No," functions as a binary classification task. Consequently, greedy search is the most suitable decoding strategy, rendering sampling-based methods unjustifiable. As shown in the upper-left corner of Figure˜4, experimental results further confirm that greedy search significantly outperforms direct sampling. In fact, greedy search outperforms direct sampling on the vast majority of tasks[42]. Additional comparisons of decoding strategies and their effects on prediction outcomes are provided in Appendix˜D. However, many contrastive decoding methods report performance improvements using sampling strategies, necessitating a closer examination of these claims.
We revisit the adaptive plausibility constraint introduced in Section˜3 and formally defined in Equation˜2. This constraint ensures that when the model exhibits high confidence in its outputs corresponding to the original input, the candidate pool is refined to retain only high-probability tokens. By incorporating this mechanism into contrastive decoding methods, it aims to mitigate adverse effects by preventing the generation of implausible tokens, while safeguarding the coherence and quality of the generated content.
Figure˜4 provides a detailed explanation of why the adaptive plausibility constraint alone significantly improves performance when using the sampling strategy. For the question "Is there a cell phone in the image?", LLaVA-v1.5-7B model generates the correct output distribution: Yes: 8.8% and No: 91.2%. Using a greedy search strategy, the model consistently produces the correct answer, "No." However, when employing a sampling strategy, there is an 8.8% chance that the model generates the incorrect answer, "Yes."
Among the excluded candidates is the option "Yes." Consequently, the sampling strategy reduces to a greedy search, ensuring a 100% probability of correctly answering "No."
Consequently, the adaptive plausibility constraint greatly limits the pool of candidate options, transforming the sampling strategy into a predominantly greedy search. As demonstrated in Figure˜4, MLLMs exhibit markedly superior performance on the POPE Benchmark under a greedy strategy, underscoring the constraint’s pivotal contribution to performance gains.
4.3 Insights
In Section˜4.1, we show that when using greedy search as the decoding strategy, contrastive decoding methods modify the model’s predictions in a unidirectional manner, shifting the output distribution toward Yes. As a result, performance improvements primarily depend on whether the model’s original output distribution was biased toward No.
In Section˜4.2, we demonstrate that when direct sampling is used as the decoding strategy, the adaptive plausibility constraint effectively reduces it to greedy search, serving as a key driver of the observed performance gains.
5 Spurious Improvement Methods
In this section, we propose a series of spurious improvement methods based on the two fundamental reasons for performance improvement discussed in Section˜4.3. Although these methods are entirely unrelated to hallucination mitigation, they yield experimental results comparable to contrastive decoding techniques. This evidence suggests that while contrastive decoding enhances performance on POPE Benchmark, it does not address hallucinations.
5.1 Forced Distribution Adjustment
For the first misleading factor in performance improvement, which involves modifying model predictions in a single direction to bias the output distribution toward "Yes," we introduce two pseudo-performance enhancement methods: Prompt-Based Adjustment and Output Layer Modification. The implementation of these algorithms is detailed in Figure˜5.
Prompt-Based Adjustment modifies the input side of the model by appending an additional prompt, "Answer Yes if possible," after the user’s instruction. This extra input biases the model’s output distribution toward "Yes." Output Layer Modification refers to adjustments made at the output stage of the model. After generating its initial prediction, the model evaluates the probabilities of "Yes" and "No." If their difference is small, i.e.,
the prediction is forcibly set to "Yes." Here, controls how close the probabilities must be to trigger this modification. This adjustment increases the likelihood of the model predicting "Yes."
Experimental Settings. We selected QwenVL-7B, LLaVA-v1.5-7B, and LLaVA-v1.5-13B as the backbone MLLMs. For decoding, we employed a greedy search strategy. Regarding the critical POPE benchmark, our experiments were conducted on COCO dataset, where the raw output distribution of MLLMs tends to be biased toward "No." As a result, all contrastive decoding methods exhibited notable improvements on this dataset. In addition, we conducted experiments on several other benchmarks that have not been evaluated by many contrastive decoding methods, including MME[43], CHAIR[44], NoCaps[45] and LLaVA-Bench[31], covering both discriminative and generative tasks.
| Category | Method | LLaVA-v1.5-7B | LLaVA-v1.5-13B | QwenVL-Chat-7B | |||
| Accuracy | Yes (%) | Accuracy | Yes (%) | Accuracy | Yes (%) | ||
| Random | Greedy | 87.1 0.0 | 39.2 0.0 | 86.7 0.0 | 38.7 0.0 | 85.9 0.0 | 37.8 0.0 |
| VCD | 88.6 1.5 | 46.4 7.2 | 89.2 2.5 | 44.4 5.7 | 87.7 1.8 | 40.6 2.8 | |
| SID | 87.9 0.8 | 42.4 3.2 | 87.2 0.5 | 42.5 3.8 | 86.5 0.6 | 39.9 2.1 | |
| PBA | 87.6 0.5 | 40.2 1.0 | 90.2 3.5 | 45.7 7.0 | 87.3 1.4 | 41.5 3.7 | |
| OLM | 89.6 2.5 | 44.2 5.0 | 90.0 3.3 | 48.8 10.1 | 88.2 2.3 | 43.8 6.0 | |
| Popular | Greedy | 85.8 0.0 | 40.4 0.0 | 86.0 0.0 | 39.4 0.0 | 85.6 0.0 | 38.2 0.0 |
| VCD | 86.2 0.4 | 48.8 8.4 | 87.3 1.3 | 46.3 6.9 | 87.1 1.5 | 41.2 3.0 | |
| SID | 85.1 0.7 | 45.1 4.7 | 85.1 0.9 | 44.6 5.2 | 85.3 0.3 | 39.8 1.6 | |
| PBA | 86.2 0.4 | 41.6 1.2 | 88.4 2.4 | 47.5 8.1 | 86.8 1.2 | 42.3 4.1 | |
| OLM | 87.3 1.5 | 46.5 6.1 | 88.6 2.6 | 50.2 10.8 | 87.4 1.8 | 44.8 6.6 | |
| Adversarial | Greedy | 83.6 0.0 | 42.6 0.0 | 84.3 0.0 | 41.0 0.0 | 84.0 0.0 | 39.7 0.0 |
| VCD | 81.9 1.7 | 53.1 10.5 | 83.8 0.5 | 49.7 8.7 | 84.5 0.5 | 43.7 4.0 | |
| SID | 82.3 1.3 | 47.9 5.3 | 82.9 1.4 | 46.9 5.9 | 83.2 0.8 | 42.5 2.8 | |
| PBA | 83.7 0.1 | 44.0 1.4 | 84.5 0.2 | 51.3 10.3 | 84.1 0.1 | 45.2 5.5 | |
| OLM | 83.6 0.0 | 50.1 7.5 | 83.9 0.4 | 54.9 13.9 | 84.8 0.8 | 48.4 8.7 | |
Results and Analysis on POPE. The experimental results, presented in Table˜3, show that PBA achieves an even greater performance improvement than SID, while OLM surpasses VCD. Prediction accuracy increases as the output distribution approaches balance. However, it is worth noting that on the Adversarial subset, when the distribution shifts beyond balance and biases toward "Yes," accuracy begins to decline, aligning with the conclusion in Section˜4.3. Although PBA and OLM are not designed for hallucination mitigation, they produce results similar to contrastive decoding methods, suggesting that contrastive decoding does not effectively address hallucinations. For the experimental results on the AOKVQA and GQA datasets, please refer to Appendix˜E.
| Method | MME-P | MME-C | CHAIR-S | CHAIR-I | Nocaps | LLaVA-Bench |
| Greedy | 1475.1 00.0 | 349.6 00.0 | 21.6 0.0 | 7.2 0.0 | 83.8 0.0 | 65.7 0.0 |
| VCD | 1389.9 85.2 | 294.6 55.0 | 25.4 3.8 | 8.1 0.9 | 81.2 2.6 | 64.6 1.1 |
| SID | 1396.4 78.7 | 304.1 45.5 | 24.9 3.3 | 7.5 0.3 | 81.9 1.9 | 64.9 0.8 |
Results and Analysis on Other Benchmarks. As shown in Table˜4, for the discriminative task MME, contrastive decoding methods did not yield any performance gains, as the model’s output did not exhibit an initial distributional bias. In fact, it led to a decline in performance. Similarly, in generative tasks such as CHAIR, Nocaps, and LLaVA-Bench, contrastive decoding offered no improvements. This clearly demonstrates that contrastive decoding can only enhance performance in discriminative tasks where the initial output distribution is skewed towards "No." However, such improvements do not fundamentally address the issue of object hallucination.
5.2 Standalone Application of the Constraint.
Experimental Settings. We utilize QwenVL-7B, LLaVA-v1.5-7B, and LLaVA-v1.5-13B as our foundational MLLMs, employing a sampling decoding strategy. Regarding the critical POPE benchmark, our experiments are conducted on the GQA dataset, where the original output distribution of MLLMs is relatively balanced. Consequently, the modification introduced by contrastive decoding methods, which shifts the output distribution towards "Yes," does not introduce a positive bias. However, since the adaptive plausibility constraint converts the sampling strategy into a greedy search, the model’s performance still improves.
| Category | Method | LLaVA-v1.5-7B | LLaVA-v1.5-13B | QwenVL-Chat-7B | |||
| Accuracy | Yes (%) | Accuracy | Yes (%) | Accuracy | Yes (%) | ||
| Random | sample | 83.8 0.0 | 45.6 0.0 | 84.6 0.0 | 45.9 0.0 | 81.5 0.0 | 41.1 0.0 |
| VCD | 86.6 2.8 | 52.5 6.9 | 86.7 2.1 | 49.5 3.6 | 83.8 2.3 | 44.0 2.9 | |
| ICD | 85.2 1.4 | 47.0 1.4 | 85.8 1.2 | 44.9 1.0 | 82.5 1.0 | 42.0 0.9 | |
| SID | 84.9 1.1 | 49.1 3.5 | 86.0 1.4 | 49.8 3.9 | 82.9 1.4 | 43.5 2.4 | |
| sample† | 85.4 1.6 | 45.1 0.5 | 86.1 1.5 | 45.3 0.6 | 83.0 1.5 | 41.8 0.7 | |
| Popular | sample | 77.3 0.0 | 52.1 0.0 | 80.6 0.0 | 49.9 0.0 | 76.8 0.0 | 46.1 0.0 |
| VCD | 78.7 1.4 | 59.4 7.3 | 82.9 2.3 | 52.4 2.5 | 78.2 1.4 | 49.4 3.3 | |
| ICD | 78.1 0.8 | 54.0 1.9 | 81.5 0.9 | 49.3 0.6 | 77.5 0.7 | 47.2 1.1 | |
| SID | 78.4 1.1 | 53.7 1.6 | 82.5 1.9 | 53.3 3.4 | 77.9 1.1 | 48.0 1.9 | |
| sample† | 78.6 1.3 | 52.0 0.1 | 81.8 1.2 | 49.6 0.3 | 78.1 1.3 | 46.8 0.7 | |
| Adversarial | sample | 75.1 0.0 | 54.1 0.0 | 78.2 0.0 | 53.2 0.0 | 76.4 0.0 | 45.5 0.0 |
| VCD | 76.4 1.3 | 62.5 8.4 | 80.3 2.1 | 57.0 3.8 | 78.6 2.2 | 49.2 3.7 | |
| ICD | 75.8 0.7 | 54.2 0.1 | 79.2 1.0 | 52.8 0.4 | 76.8 0.4 | 46.0 0.5 | |
| SID | 76.3 1.2 | 57.5 3.4 | 78.7 0.5 | 57.5 4.3 | 77.2 0.8 | 47.5 2.0 | |
| sample† | 76.3 1.2 | 54.2 0.1 | 79.5 1.3 | 53.1 0.1 | 77.9 1.5 | 46.2 0.7 | |
Results and Analysis on POPE. The experimental results in Table˜5 show that when MLLMs adopt sampling as the decoding strategy, applying the adaptive plausibility constraint alone yields an approximate 2.5% performance improvement, effectively validating the conclusion in Section˜4.3. Notably, since the adaptive plausibility constraint is entirely unrelated to hallucination mitigation yet achieves performance on par with various contrastive decoding methods, this strongly suggests that contrastive decoding methods do not actually mitigate hallucinations. For the experimental results on the AOKVQA and COCO datasets, please refer to Appendix˜F.
Results and Analysis on Other Benchmarks. As shown in Table˜6, across all other benchmarks, employing the adaptive plausibility constraint alone under direct sampling yields the most significant performance gains. This suggests that the main reason behind the performance improvements observed with contrastive decoding lies in the fact that the constraint reduces direct sampling to greedy search. Crucially, this mechanism is unrelated to the original goal of mitigating hallucination.
| Method | MME-P | MME-C | CHAIR-S | CHAIR-I | Nocaps | LLaVA-Bench |
| Sample | 1282.1 0.0 | 286.7 0.0 | 23.2 0.0 | 7.7 0.0 | 81.4 0.0 | 64.3 0.0 |
| Sample† | 1392.4 110.3 | 302.6 15.9 | 22.4 0.8 | 7.4 0.3 | 83.1 1.7 | 65.3 1.0 |
| VCD | 1361.7 79.6 | 278.6 8.1 | 25.8 2.6 | 8.3 0.6 | 80.9 0.5 | 64.2 0.1 |
| SID | 1364.6 82.5 | 284.1 2.6 | 25.2 2.0 | 7.9 0.2 | 81.2 0.2 | 64.5 0.2 |
6 Discussion on Hallucination Mitigation
Based on the insights from Sections˜4 and 5, we propose some new criteria for evaluating object hallucination mitigation in MLLMs.
Impact of Decoding Strategies. When evaluating on POPE, it is essential to account for the substantial influence of different decoding strategies on model performance. Notably, greedy search consistently outperforms sampling-based methods such as direct sampling and nucleus sampling. If a hallucination mitigation method involves modifications to the sampling module, careful consideration must be given to whether these changes affect the core properties of the decoding strategy.
Balancing Correction and Preservation. An effective hallucination mitigation method must strike a balance: it should correct incorrect answers while preserving correct ones. However, as shown in Figure˜2, some flawed approaches, despite fixing many errors, also introduce unnecessary modifications to originally correct responses. This behavior resembles mere answer editing rather than genuine hallucination mitigation. To enhance evaluation rigor, future studies should explicitly report instances where correct responses are mistakenly altered, providing a clearer measure of a method’s true effectiveness.
7 Conclusion
This study demonstrates that the performance improvements of contrastive decoding on the POPE benchmark largely stem from two misleading factors: (1) a unidirectional shift in the model’s output distribution, which biases it toward generating "Yes" responses, artificially balancing the distribution in certain datasets, and (2) the adaptive plausibility constraint, which reduces sampling decoding to greedy search. By comparing experimental results from spurious improvement methods and contrastive decoding, we confirm that while contrastive decoding enhances performance, it ultimately fails to mitigate hallucinations.
Impact Statement
Acknowledgements
This work is supported by the National Natural Science Foundation of China under Grant 62176246. This work is also supported by Anhui Province Key Research and Development Plan (202304a05020045), Anhui Province Natural Science Foundation (2208085UD17) and National Natural Science Foundation of China under Grant 62406098.
Appendix A Contrastive Decoding for Hallucinations